# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v1

@@ Abstract
Mechanically interlocked polymers (MIPs) are a novel class of polymer structures in which the components are connected by mechanical bonds instead of covalent bonds. We measure the single-molecule rheological properties of polyrotaxanes, daisy chains, and polycatenanes under steady shear and steady uniaxial extension using coarse-grained Brownian dynamics simulations with hydrodynamic interactions. We obtain key rheological features, including tumbling dynamics, molecular extension, stress, and viscosity. By systematically varying structural features, we demonstrate how MIP topology governs flow response. Compared to linear polymers, all three MIP architectures exhibit enhanced tumbling in shear flow and lower normal stress differences in extensional flow. While polyrotaxanes show higher shear and extensional viscosities, polycatenanes and daisy chains have lower viscosities. In extensional flow, polyrotaxanes and polycatenanes extend earlier than linear polymers. We find that mechanical bonds suppress shear thinning and alter the coil-stretch transition observed in linear polymers. These effects arise from the mechanically bonded rings in MIPs, which expand the polymer profile in gradient direction and increase backbone stiffness due to ring-backbone repulsions. This study provides key insights into MIP flow properties, providing the foundation for their systematic development in engineering applications.
@@ I Introduction
Mechanically interlocked polymers (MIPs) are an emerging class of polymer architectures distinguished by components linked through mechanical entanglements rather than covalent bonds. These unique structural features impart MIPs with increased conformational freedom, setting them apart from typical polymers Hart et al. (2021) and leading to unique propertiesSluysmans and Stoddart (2019); Liu et al. (2022) such as shape memory, ductile glassy states, large strain-at-break, scratch resistance, and stimuli-responsive mechanical properties. Theoretical and computational studies have further indicated that they have large loss modulus, low activation energy for viscous flow, superior energy-damping, and excellent toughnessAndo and Ito (2025); Wu et al. (2017); Hart et al. (2021).
@@
The schematic representation in Fig. 1 shows different classes of MIPs in comparison with a linear polymer (Fig. 1a). Polyrotaxanes (Fig. 1b) consist of multiple ring-shaped molecules threaded onto a linear backbone, held in place by bulky end caps called stoppersHarrison and Harrison (1967); Wenz and Keller (1992). Daisy chains (Fig. 1c) are made from covalently-bonded rings and linear sections interlocked in seriesAshton et al. (1998); Wenz et al. (2006). Polycatenanes (Fig. 1d) are polymers made of interlocked rings, like a chainSchill and Lüttringhaus (1964); Frisch and Wasserman (1961); Weidmann et al. (1999); Wu et al. (2017). Although mechanically interlinking molecules (catenation) is a synthetic challenge Segawa et al. (2019); Leigh et al. (2014), recent advances have enabled the production of MIPs with moderate molecular weights. The largest polyrotaxane synthesized to date has weight-average molecular weight $M_{w}$ = 119,000 g/mol with 86 threaded rings Araki et al. (2005) while the largest daisy chain polymer has 11 repeat units and number-average molecular weight $M_{n}\approx$ 25,000 g/mol Cai et al. (2020). The longest linear and branched polycatenanes ever built have 26 and 130 rings in their structures with $M_{n}\approx$ 40,000 g/mol and $M_{n}\approx$ 200,000 g/mol, respectively Wu et al. (2017). Nano-scale polycatenanes comprising up to 22 interlocked rings have been synthesized through the self-assembly of supramolecular toroidal building blocks Datta et al. (2020).
@@
Despite advances in synthesis, the relationship between mechanical bonds and flow properties is largely unexplored. Such properties need to be measured to develop MIPs for practical applications. Computational techniques present an effective method to link the molecular structure, dynamics, and macroscopic behavior of polymersGartner III and Jayaraman (2019); Li et al. (2013); Gedde et al. (2021); Seyedi et al. (2020); Kremer and Grest (1990). In the realm of MIPs, computational research has quantified the structural and dynamic properties of doubly interlocked rings (2-catenanes) in equilibrium conditions Rauscher et al. (2020); Bohn et al. (2010), in shear Li et al. (2025), and in glassy states Li et al. (2023). Simulation studies have explored the equilibrium behavior of longer polycatenanes (up to 300 rings), isolated Tubiana et al. (2022); Chiarantoni and Micheletti (2022); Rauscher et al. (2018); Dehaghani et al. (2020); Pakula and Jeszka (1999); Li et al. (2021) and in a melt state Rauscher et al. (2020, 2020); Hagita et al. (2021), uncovering characteristics that blend features of ring and linear polymers. In addition, computational investigations have probed the structure and dynamics of catenanes under confinement Chiarantoni and Micheletti (2023) and during pore translocation Caraglio et al. (2017); Suma and Micheletti (2017) and rotaxanes have been analyzed in equilibrium Li et al. (2021); Yasuda et al. (2019). These simulations have highlighted the unique behavior of MIPs compared to conventional polymers, demonstrating how structural features, such as ring size and the number of mechanical bonds, influence properties like chain stiffness, $\theta$-temperature, and relaxation dynamics. Importantly, these studies show that MIPs exhibit emergent properties not observed in typical linear and cyclic polymer architectures, highlighting the crucial role of topology in governing their behavior.
@@
Here we simulate MIPs with coarse-grained Brownian dynamics, an approach that has been effective at characterizing the rheology of many polymer morphologies, like linear, combs, bottlebrushes, and ring polymersMai et al. (2018); Lyulin et al. (1999); Young et al. (2019); Dutta and Sing (2024). We report rheological properties for polyrotaxanes, daisy chains, and polycatenanes in both steady shear and steady extensional flow, a critically uncharacterized domain for MIPs. Our results are a first step to understanding how MIPs will behave during processing as extensional flows are commonly encountered in blow molding, fiber spinning, film blowing, and foaming Bayer et al. (1989); Sentmanat (2004); Starỳ et al. (2015) and shear flows appear in injection molding and extrusion Martin et al. (2015). We show that all three MIP types show enhanced tumbling in shear flow and lower normal stress differences in extensional flow. Also, while polyrotaxanes show higher viscosity both in shear and extension, polycatenanes and daisy chains exhibit lower viscosities. In addition, polyrotaxanes and polycatenanes have an earlier onset of extension than linear polymers under extensional flow.
@@
Figure 1: Schematic representation of polymers: (a) linear polymer, (b) polyrotaxane, (c) daisy chain, and (d) polycatenane. Yellow and blue beads are equivalent, the colors are chosen to highlight the interlocked architecture. Green beads are large end caps to prevent dethreading in polyrotaxanes and daisy chains.
@@ II Methods
We simulated single polymers in the dilute solution limit with a coarse-grained bead-based model. Polymers are composed of $N$ beads that interact through bonded and non-bonded potentials. The pairwise non-bonded interaction potential is defined by the Weeks–Chandler–Anderson potential Weeks et al. (1971)
@@
where $r_{ij}=\left|\mathbf{r}_{j}-\mathbf{r}_{i}\right|$ is the distance between particles $i$ and $j$, $r_{c}$ is the cutoff distance, $U_{c}=4\epsilon\left[\left(\frac{\sigma}{r_{c}}\right)^{12}-\left(\frac{\sigma}{r_{c}}\right)^{6}\right]$ is the Lennard-Jones energy at the cutoff distance, and $\sigma$ and $\epsilon$ are the Lennard-Jones length and energy parameters, respectively. To prevent the rings from dethreading, the non-bonded interactions between capping beads and ring beads are scaled by a factor of $10^{5}$ and $10^{10}$ in daisy chains and polyrotaxanes, respectively. Bonds between beads are modeled with finitely extensible nonlinear elastic (FENE) springs Kremer and Grest (1990):
@@
where $k_{s}=30\varepsilon/\sigma^{2}$ is the spring strength and $r_{\mathrm{max}}=1.5\sigma$ is the maximum bond extension. This model and parameterization has been used to model MIPs Rauscher et al. (2020); Becerra et al. (2024). The positions of beads evolve according to the following stochastic differential equation incorporating hydrodynamic interactions Ermak and McCammon (1978); Ottinger (1996):
@@
where $\mathbf{r}$ is a vector of particle positions with 3N elements, $\mathbf{v}$ is the solvent flow velocity at the particle positions, $\mathbf{f}$ is a vector of forces acting on the particles, $\mathbf{D}$ is the diffusivity matrix, $\mathbf{B}=\sqrt{\mathbf{D}}$, and $\mathbf{w}$ is a vector containing independent Wiener processes with zero mean and variance $dt$. In Eq. 3, the deterministic drift is calculated by the term within the square brackets, while the other term accounts for the Brownian motion of the particles due to random collisions with solvent molecules. The force vector is $\mathbf{f}=-\nabla U$, where $U=U_{\text{WCA}}+U_{\text{FENE}}$ is the total intramolecular potential. The matrix $\mathbf{D}$ incorporates hydrodynamic interactions between the particles. Each $3\times 3$ block of the matrix $\mathbf{D}$ represents hydrodynamic interactions between particles $i$ and $j$ and is given by
@@
is the Rotne–Prager–Yamakawa tensor Rotne and Prager (1969); Yamakawa (1970). To remove the overall translational motion of the molecule, the solvent velocity is centered around the center-of-mass of the polymer $\mathbf{r}_{\mathrm{COM}}=[x_{\mathrm{COM}},y_{\mathrm{COM}},z_{\mathrm{COM}}]^{\top}$ Miao et al. (2017). Thereby, the solvent velocity $\mathbf{v}_{i}$ for bead $i$ with spatial coordinate $\mathbf{r}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ is zero at equilibrium; $[\dot{\gamma}(y_{i}-y_{\mathrm{COM}}),0,0]^{\top}$ in shear flow, where $\dot{\gamma}$ is the shear rate; and $[\dot{\varepsilon}(x_{i}-x_{\mathrm{COM}}),-\dot{\varepsilon}(y_{i}-y_{\mathrm{COM}})/2,-\dot{\varepsilon}(z_{i}-z_{\mathrm{COM}})/2]^{\top}$ in uniaxial extensional flow, where $\dot{\varepsilon}$ is the extension rate. Furthermore, the Weissenberg number (Wi) measures the flow strength and is defined as $\dot{\gamma}\tau_{R}$ for shear flow and $\dot{\varepsilon}\tau_{R}$ for uniaxial extensional flow, where $\tau_{R}$ is the longest relaxation time of the polymer as calculated from $\tau_{R}=R_{g}^{2}/6D_{\mathrm{COM}}$ where the center-of-mass diffusivity was calculated from the equilibrium conformations using the Kirkwood formalism Dünweg et al. (2002). Relaxation time results are summarized in tables A.1 to A.4. The variables are nondimensionalized using $a$ as the unit of length, $k_{B}T$ as the unit of energy, and the diffusion time of a bead $\tau=\zeta a^{2}/{k_{B}T}$ as the unit of time. Eq. 3 is integrated numerically using a semi-implicit Euler scheme Jendrejack et al. (2000, 2002) to allow the use of larger time steps compared to explicit methods. The diffusivity matrix $\mathbf{D}$ and the Brownian term are treated explicitly and the remaining terms are treated implicitly.
The following set of nonlinear algebraic equations in $\mathbf{r}^{*}$ are obtained using the semi-implicit algorithm:
@@
Equation 6 is solved iteratively by a Jacobian-free Newton–Krylov method Knoll and Keyes (2004) through the NITSOL software package Pernice and Walker (1998) without using a preconditioner. This approach has recently been used to simulate bottlebrush polymers Dutta and Sing (2024). We use OpenMP shared memory parallelism with 1-4 CPUs per simulation (depending on system size) to more efficiently calculate forces, perform the Cholesky decomposition of the $\mathbf{D^{*}}$ matrix, and integrate the equation of motion. Initial configurations were generated with equilibrium bond lengths and energy minimization was carried out prior to production simulations. All properties were calculated by averaging over 10 to 20 trajectories of at least 100 relaxation times long ($10^{6}-10^{8}$ time steps) with time steps ranging from $10^{-2}$ to $10^{-4}$ depending on the strain rate. The fractional extension is defined as $\langle X_{bb}\rangle/L$ where $X_{bb}$ is the difference between minimum and maximum coordinates in the $x$ direction and $L$ is the maximum contour length, calculated as $L=3a(N_{\mathrm{eff}}-1)$. $N_{\mathrm{eff}}$ is effective number of beads which is determined differently depending on the polymer type. For linear polymers, $N_{\mathrm{eff}}=N$ is the total number of beads. In the case of polyrotaxanes, $N_{\mathrm{eff}}=N_{bb}$ is the number of beads in the linear backbone. For polycatenanes, $N_{\mathrm{eff}}=N/2$ is calculated as the total number of beads divided by 2. This adjustment accounts for the contour length of ring polymers, which is considered to be half the contour length of the corresponding linear polymer Zhou et al. (2019). For daisy chains, $L$ is calculated separately for the rings and the linear segments and then summed.
@@
Under steady uniaxial extension, we calculate the fractional extension $\langle X_{bb}\rangle/L$, extensional viscosity $\eta_{e}$ and first normal stress difference $N_{1}$ across a range of Wi values. Under steady shear flow, we calculate the fractional extension in $x$, $y$ and $z$ directions ($\langle X_{bb}\rangle/L$, $\langle Y_{bb}\rangle/L$, $\langle Z_{bb}\rangle/L$), first normal stress coefficient ($\psi_{1}$), shear viscosity $\eta$, orientation angle $\theta$, and tumbling frequency ($\omega\tau_{R}$) over a range of Wi values. We use the radius of gyration along the flow and gradient directions to calculate the orientation angle $\theta$ to measure the alignment of the chains in the $xy$ plane using the following equation Schroeder et al. (2005).
$\tan(2\theta)=\frac{2R_{g,xy}}{R_{g,xx}-R_{g,yy}}$ (7)
To quantify and analyze the tumbling dynamics of the polymers under shear flow, we track and analyze the evolution of the angle between the end-to-end vector of the polymer chains and the $x$-axis, similar to the work of Dalal et. al. Saha Dalal et al. (2012). We monitor the angle as it progresses through the four quadrants of a full $360^{\circ}$ rotation cycle and record the time taken for the chains to complete each full cycle (i.e., one complete rotation through all four quadrants) throughout the simulation and use the mean value as the tumbling time. We calculate tumbling frequency as inverse tumbling time, normalized by the relaxation time ($\tau_{R}$).

### v2

@@ Abstract
Mechanically interlocked polymers (MIPs) are a class of polymer structures in which the components are connected by mechanical bonds instead of covalent bonds. We measure the single-molecule rheological properties of polyrotaxanes, daisy chains, and polycatenanes under steady shear and steady uniaxial extension using coarse-grained Brownian dynamics simulations with hydrodynamic interactions. We obtain key rheological features, including tumbling dynamics, molecular extension, stress, and viscosity. By systematically varying structural features, we demonstrate how MIP topology governs flow response. Compared to linear polymers, all three MIP architectures exhibit enhanced tumbling in shear flow, weaker shear thinning, and lower normal stress differences in extensional flow. While polyrotaxanes show higher shear and extensional viscosities, polycatenanes and daisy chains have lower viscosities. In extensional and shear flows, MIPs typically extend more in the flow direction and at weaker flow strength than linear polymers. These effects arise from the mechanically bonded rings in MIPs, which expand the polymer profile in gradient direction and increase backbone rigidity due to ring-backbone repulsions. This study provides key insights into MIP flow properties, providing the foundation for their systematic development in engineering applications.
@@ I Introduction
Mechanically interlocked polymers (MIPs) are a class of polymer architectures distinguished by components linked with mechanical entanglements rather than covalent bonds. The schematic representation in Fig. 1 shows different classes of MIPs in comparison with a typical linear polymer (Fig. 1a). Polyrotaxanes (Fig. 1b) consist of multiple ring-shaped molecules threaded onto a linear backbone, held in place by bulky end caps called stoppersHarrison and Harrison (1967); Wenz and Keller (1992). Daisy chains (Fig. 1c) are made from covalently-bonded rings and linear sections interlocked in seriesAshton et al. (1998); Wenz et al. (2006). Polycatenanes (Fig. 1d) are polymers made of interlocked rings, like a chainSchill and Lüttringhaus (1964); Frisch and Wasserman (1961); Weidmann et al. (1999); Wu et al. (2017). These mechanical bonds impart MIPs with increased conformational freedom, setting them apart from typical polymers Hart et al. (2021) and leading to unique propertiesSluysmans and Stoddart (2019); Liu et al. (2022) such as shape memory, ductile glassy states, large strain-at-break, scratch resistance, and stimuli-responsive mechanical properties Hart et al. (2021); Zhou et al. (2024); Ando and Ito (2025). From theoretical studies it is anticipated that MIPs will show large loss modulus, low activation energy for viscous flow, and rapid stress relaxation Weidmann et al. (1999), potentially acting as superior energy-damping materials with excellent toughness Wu et al. (2017). In addition, equilibrium and strain simulations of rotaxanes and rotaxane-based materials, like slide-ring gels and tendomer gels, have demonstrated exceptional elasticity and yield stresses surpassing conventional cross-linked polymer gels Li et al. (2021); Yasuda et al. (2019); Chen et al. (2023); Tanahashi and Koga (2021); Uehara et al. (2022); Yasuda et al. (2020); Müller et al. (2019, 2021, 2022). These examples demonstrate the unique properties imparted by mechanical bonds.
@@
Although mechanically interlinking molecules is a synthetic challenge Segawa et al. (2019); Leigh et al. (2014), recent advances have enabled the production of MIPs with moderate molecular weights. The largest polyrotaxane synthesized to date has weight-average molecular weight $M_{w}$ = 119,000 g/mol with 86 threaded rings Araki et al. (2005) while the largest daisy chain polymer has 11 repeat units and number-average molecular weight $M_{n}\approx$ 25,000 g/mol Cai et al. (2020). Ultra-high-molecular-weight polycatenanes have recently been synthesized, consisting of up to 2,800 rings with a molecular weight of 4.5 MDa Wang et al. (2025). Nano-scale polycatenanes comprising up to 22 interlocked rings have been synthesized through the self-assembly of supramolecular toroidal building blocks Datta et al. (2020).
@@
Despite advances in synthesis, the relationship between mechanical bonds and flow properties is largely unexplored. Computational techniques present an effective method to link the molecular structure, dynamics, and macroscopic behavior of polymersGartner III and Jayaraman (2019); Li et al. (2013); Gedde et al. (2021); Seyedi et al. (2020); Kremer and Grest (1990). In the realm of MIPs, simulations have quantified the structural and dynamic properties of doubly interlocked rings (2-catenanes) in equilibrium conditions Rauscher et al. (2020); Bohn et al. (2010), in shear Li et al. (2025), and in glassy states Li et al. (2023). Simulation studies have explored the equilibrium behavior of longer polycatenanes (up to 300 rings), isolated Tubiana et al. (2022); Chiarantoni and Micheletti (2022); Rauscher et al. (2018); Dehaghani et al. (2020); Pakula and Jeszka (1999); Li et al. (2021) and in a melt state Rauscher et al. (2020, 2020); Hagita et al. (2021), uncovering characteristics that blend features of ring and linear polymers. In addition, computational investigations have probed the structure and dynamics of catenanes under confinement Chiarantoni and Micheletti (2023) and during pore translocation Caraglio et al. (2017); Suma and Micheletti (2017) and rotaxanes have been analyzed in equilibrium Li et al. (2021); Yasuda et al. (2019). These simulations have highlighted the unique behavior of MIPs compared to conventional polymers, demonstrating how structural features, such as ring size and the number of mechanical bonds, influence properties like chain stiffness Pakula and Jeszka (1999); Rauscher et al. (2020), $\theta$-temperature Dehaghani et al. (2020); Guo et al. (2023), and relaxation dynamics Pakula and Jeszka (1999); Rauscher et al. (2018). Importantly, these studies show that MIPs exhibit emergent properties not observed in typical linear and cyclic polymer architecturesPakula and Jeszka (1999), highlighting the crucial role of topology in governing their behavior.
@@
Here we simulate dilute solutions of MIPs with coarse-grained Brownian dynamics, an approach that has been effective at characterizing the rheology of many polymer morphologies, like linear Lyulin et al. (1999), combs Mai et al. (2018), bottlebrushes Dutta and Sing (2024), and ring polymers Young et al. (2019). We report rheological properties for polyrotaxanes, daisy chains, and polycatenanes in both steady shear and steady extensional flow, a critically uncharacterized domain for MIPs. We show that all three MIP types show enhanced tumbling in shear flow, weaker shear thinning, and lower normal stress differences in extensional flow. Also, while polyrotaxanes show higher viscosity both in shear and extension, polycatenanes and daisy chains exhibit lower viscosities. Polymer extension is typically greater in both shear and extensional flows, and approaches linear values as the density of mechanical bonds in MIPs decreases.
@@
Figure 1: Schematic representation of polymers: (a) linear polymer, (b) polyrotaxane, (c) daisy chain, and (d) polycatenane. Yellow and blue beads are equivalent, the colors are chosen to highlight the interlocked architecture. Green beads are end caps to prevent dethreading in polyrotaxanes and daisy chains.
@@ II Methods
We simulated single polymers in the dilute solution limit with a coarse-grained bead-based model. Using a previously established model for MIPs Rauscher et al. (2020); Becerra et al. (2024), each polymer consists of $N$ beads that interact through bonded and non-bonded potentials. The pairwise non-bonded interaction potential is the Weeks–Chandler–Anderson potential Weeks et al. (1971)
@@
where $r_{ij}=\left|\mathbf{r}_{j}-\mathbf{r}_{i}\right|$ is the distance between particles $i$ and $j$, $r_{c}=2.2449241$ is the minimum energy cutoff distance, $U_{c}=4\epsilon\left[\left(\frac{\sigma}{r_{c}}\right)^{12}-\left(\frac{\sigma}{r_{c}}\right)^{6}\right]$ is the Lennard-Jones energy at the cutoff distance, and $\sigma$ and $\epsilon$ are the Lennard-Jones length and energy parameters, respectively. To prevent polyrotaxane and daisy chain rings from dethreading, we introduced non-bonded interactions between capping beads and ring beads that are modeled without the attractive term, shift and cutoff by using $U_{\mathrm{cap}}=4\epsilon_{\mathrm{cap}}\left(\frac{\sigma}{r_{ij}}\right)^{12}$, where $\epsilon_{\mathrm{cap}}=10^{5}\epsilon$ in daisy chains and $\epsilon_{\mathrm{cap}}=10^{10}\epsilon$ in polyrotaxanes. Due to the absence of an attractive term, these $\epsilon_{\mathrm{cap}}$ values are equivalent to 2.6 and 6.8 fold increase in $\sigma$ size parameter, respectively. Bonds between beads are modeled with finitely extensible nonlinear elastic (FENE) springs Kremer and Grest (1990):
@@
where $k_{s}=30\varepsilon/\sigma^{2}$ is the spring strength and $r_{\mathrm{max}}=1.5\sigma$ is the maximum bond extension.
The bead positions evolve according to the following stochastic differential equation incorporating hydrodynamic interactions and implicit solvent Ermak and McCammon (1978); Ottinger (1996):
@@
where $\mathbf{r}$ is a vector of particle positions with $3N$ elements, $\mathbf{v}$ is the solvent flow velocity at the particle positions, $\mathbf{f}$ is a vector of forces acting on the particles, $\mathbf{D}$ is the diffusivity matrix, $\mathbf{B}=\sqrt{\mathbf{D}}$, and $\mathbf{w}$ is a vector containing independent Wiener processes with zero mean and variance $dt$. In Eq. 3, the term in the brackets is the deterministic drift, while the other term accounts for the Brownian motion of the particles due to random collisions with solvent molecules. The force vector is $\mathbf{f}=-\nabla U$, where $U=U_{\text{WCA}}+U_{\text{FENE}}+U_{\mathrm{cap}}$ is the total intramolecular potential. The matrix $\mathbf{D}$ incorporates hydrodynamic interactions between the particles. Each $3\times 3$ block of the matrix $\mathbf{D}$ represents hydrodynamic interactions between particles $i$ and $j$ and is given by
@@
is the Rotne–Prager–Yamakawa tensor Rotne and Prager (1969); Yamakawa (1970). To remove the overall translational motion of the molecule, the solvent velocity is centered around the center of mass of the polymer $\mathbf{r}_{\mathrm{COM}}=[x_{\mathrm{COM}},y_{\mathrm{COM}},z_{\mathrm{COM}}]^{\top}$ Miao et al. (2017). The solvent velocity $\mathbf{v}_{i}$ for bead $i$ with spatial coordinate $\mathbf{r}_{i}=[x_{i},y_{i},z_{i}]^{\top}$ is zero at equilibrium; $\dot{\gamma}(y_{i}-y_{\mathrm{COM}}),0,0]^{\top}$ in shear flow, where $\dot{\gamma}$ is the shear rate; and $\dot{\varepsilon}(x_{i}-x_{\mathrm{COM}}),-\dot{\varepsilon}(y_{i}-y_{\mathrm{COM}})/2,-\dot{\varepsilon}(z_{i}-z_{\mathrm{COM}})/2]^{\top}$ in uniaxial extensional flow, where $\dot{\varepsilon}$ is the extension rate. Furthermore, the Weissenberg number (Wi) measures the flow strength and is defined as $\dot{\gamma}\tau_{R}$ for shear flow and $\dot{\varepsilon}\tau_{R}$ for uniaxial extensional flow, where $\tau_{R}$ is the longest relaxation time of the polymer. Relaxation time is $\tau_{R}=R_{g}^{2}/6D_{\mathrm{COM}}$, where the center-of-mass diffusivity was calculated from the equilibrium conformations using the Kirkwood formalism Dünweg et al. (2002). Relaxation time results are summarized in SI Tables A.1 to A.4. The variables are nondimensionalized using $a$ as the unit of length, $k_{B}T$ as the unit of energy, and the diffusion time of a bead $\tau=\zeta a^{2}/{k_{B}T}$ as the unit of time. Eq. 3 is integrated numerically using a semi-implicit Euler scheme Jendrejack et al. (2000, 2002) to allow the use of larger time steps compared to explicit methods. The diffusivity matrix $\mathbf{D}$ and the Brownian term are treated explicitly and the remaining terms are treated implicitly. The following set of nonlinear algebraic equations in $\mathbf{r}^{*}$ are obtained using the semi-implicit algorithm:
@@
Equation 6 is solved iteratively by a Jacobian-free Newton–Krylov method Knoll and Keyes (2004) through the NITSOL software package Pernice and Walker (1998) without using a preconditioner. This approach has recently been used to simulate bottlebrush polymers Dutta and Sing (2024). We use OpenMP shared memory parallelism with 1-4 CPUs per simulation (depending on system size) to more efficiently calculate forces, perform the Cholesky decomposition of the $\mathbf{D^{*}}$ matrix, and integrate the equation of motion. Initial configurations were generated with equilibrium bond lengths and energy minimization was carried out prior to production simulations. All properties were calculated by averaging over 10 to 20 trajectories of at least 100 relaxation times long ($10^{6}-10^{8}$ time steps) with time steps ranging from $10^{-2}$ to $10^{-4}$ depending on the strain rate. Error bars in the graphs represent the standard error of the mean. We calculated power-law exponents and their error using parametric bootstrap (Monte Carlo resampling) Efron (1992). The fractional extension in the flow direction is $\langle X_{bb}\rangle/L$, where $X_{bb}$ is the difference between minimum and maximum coordinates in the $x$ direction and $L$ is the maximum contour length. The maximum contour length is L=3a(N1)L=3a(N-1) and $L=3a(N_{bb}-1)$ for linear polymers and polyrotaxanes respectively, where $N$ is the total number of beads and $N_{bb}$ is the number of backbone beads. For daisy chains and polycatenanes, a pair of equal and opposite forces is applied to the end beads of the polymers in the $x$-direction and gradually increased under zero-temperature conditions. The polymer extension just before a mechanical bond fails, corresponding to a maximum in stress ($\tau^{p}_{xx}$) is recorded as the maximum contour length, $L$.
@@
Under steady uniaxial extension, we calculate the fractional extension $\langle X_{bb}\rangle/L$, extensional viscosity $\eta_{e}$ and first normal stress difference $N_{1}$ across a range of Wi values. Under steady shear flow, we calculate the fractional extension in $x$ (flow), $y$ (gradient), and $z$ (cross-flow) directions ($\langle X_{bb}\rangle/L$, $\langle Y_{bb}\rangle/L$, $\langle Z_{bb}\rangle/L$), first normal stress coefficient ($\psi_{1}$), shear viscosity $\eta$, orientation angle with respect to the flow $\theta$, and tumbling frequency ($\omega\tau_{R}$) over a range of Wi values. In shear, we track and analyze the evolution of the angle between the end-to-end vector of the polymer chains and the $x$-axis to calculate the tumbling time, similar to the work of Dalal et. al. Saha Dalal et al. (2012). We calculate tumbling frequency as inverse tumbling time, normalized by the relaxation time ($\tau_{R}$).